OpenStudy (anonymous):

whats the integral of sqrt(1+t^2) ?

OpenStudy (anonymous):

[ln[|sqrt(t^2+1)+t|]/2] + [tsqrt(t^2+1)/2

OpenStudy (anonymous):

\[\frac{\ln|\sqrt{t^2+1}+t|}{2}+\frac{t\sqrt{t^2+1}}{2}\]